OpenStudy (anonymous):

Variable y varies directly with x2, and y = 72 when x = 6. What is the value of y when x = 4? y =

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

y=k(x^2) y=72 x=6 72=k(36) What is k?

OpenStudy (anonymous):

2?

OpenStudy (dtan5457):

Yes. So k=2. What is the value of y when x = 4? y=2(x^2) x=4 y=2(4^2) y=2(16) y=?

OpenStudy (anonymous):

32?

OpenStudy (dtan5457):

yep

OpenStudy (anonymous):

Yay! so y=32?